OpenStudy (anonymous):

Solve the system by elimination. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 Solve the system by substitution. -x-y-z=-8 -4+4y+5z=7 2x+2z=4

OpenStudy (anonymous):

To solve by elimination add the first equation and second equation together which will give you only 1 equation with x in it. Then you should be able to solve. For subsitiution, solve for one variable (x would be easiest and then substitute for x into the other equations. Does that make sense?

OpenStudy (anonymous):

It does kinda make sense. Can you show me the steps so I could understand it more? I've heard what you said before but I never get it. I'm really new to these systems, and I need to know how to do them by tomorrow, because I have a test.

OpenStudy (anonymous):

And I really appreciate you replying to my question. thank you.(:

OpenStudy (jdoe0001):

timashorty have you done system of equations by elimination with only 2 variables?

OpenStudy (anonymous):

yes I have but I still struggle with that too. There's so many steps, different systems. its complicated to remember all that. And I just started algebra2 a week ago, and I'm late in class the teacher skipped alot of lessons for me, which is why I'm lost with these two equations.

OpenStudy (anonymous):

-2x + 2y + 3z = 0 -2x - y + z = -3--->(-1) -------------------- -2x + 2y + 3z = 0 2x + y - z = 3 (result of multiplying by -1) -------------------add 3y + 2z = 3 we eliminated the x out of the equation, so with the last two equations x has to also be eliminated. Remember, when doing problems like this, the same variable has to be eliminated. -2x - y + z = -3 2x + 3y + 3z = 5 --------------- 2y + 4z = 2 now we take our answers and eliminate another variable.. 3y + 2z = 3 -->(-2) 2y + 4z = 2 ----------- -6y - 4z = -6 (result of multiplying by -2) 2y + 4z = 2 -----------add -4y = -4 y = 1 now we sub 1 in for y in one of the equations that does not have an x term. 2y + 4z = 2 2(1) + 4z = 2 2 + 4z = 2 4z = 2 - 2 4z = 0 z = 0/4 = 0 now we sub 1 in for y and 0 in for z in one of the original equations 2x + 3y + 3z = 5 2x + 3(1) + 3(0) = 5 2x + 3 = 5 2x = 5 - 3 2x = 2 x = 1 check... -2x + 2y + 3z = 0 -2(1) + 2(1) + 3(0) = 0 -2 + 2 + 0 = 0 0 = 0 (correct) x = 1, y = 1, and z = 0

OpenStudy (anonymous):

Thank you so much. You helped alot!

OpenStudy (anonymous):

Can you help me with the second one two?